OpenStudy (anonymous):

Solve the equation: dx/dt - x^3 = x

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{1}{x+x^{3}}dx =\int\limits_{}^{}dt\]

OpenStudy (anonymous):

Thanks